The following table gives the winning percentages of the top five coaches in NFL history. What fraction of his games has each coach won?
Don Shula 0.676
George Halas 0.672
Tom Landry 0.606
Curly Lambeau 0.624
Paul Brown 0.667

Answer by edjones(8007)

Shula: 676/1000
=169/250 divide top and bottom by 4.
.
Halas: 672/1000
=168/250 divide by 4
=84/125 divide by 2
.
Now that you know how, I'll leave the rest up to you.
.
Ed